# Physically-based Lighting Augmentation for Robotic Manipulation

### v1

@@ Abstract
Despite advances in data augmentation, policies trained via imitation learning still struggle to generalize across environmental variations such as lighting changes. To address this, we propose the first framework that leverages physically-based inverse rendering for lighting augmentation on real-world human demonstrations. Specifically, inverse rendering decomposes the first frame in each demonstration into geometric (surface normal, depth) and material (albedo, roughness, metallic) properties, which are then used to render appearance changes under different lighting. To ensure consistent augmentation across each demonstration, we fine-tune Stable Video Diffusion on robot execution videos for temporal lighting propagation. We evaluate our framework by measuring the structural and temporal consistency of the augmented sequences, and by assessing its effectiveness in reducing the behavior cloning generalization gap (40.1%) on a 7-DoF robot across 6 lighting conditions using 720 real-world evaluations. We further showcase three downstream applications enabled by the proposed framework.
Keywords: Robotic Manipulation, Lighting Augmentation, Imitation Learning
Figure 1: Our framework enables relighting of real-world human demonstrations through inverse rendering, improving imitation learning performance under diverse lighting conditions.
1 Introduction
Imitation learning offers a promising approach to deploying various robotic manipulation tasks using expert human demonstrations [1, 2, 3]. However, as pointed out by previous research [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], a critical challenge remains: the trained policies often struggle to generalize across variations in lighting, object texture and visual distractors. Capturing variations in these individual factors in real-world datasets is resource-intensive, and this challenge is further compounded when considering their combinations, making it difficult to address with real-world data collection alone.
Image augmentation techniques are widely adopted to improve generalization across environmental variations without requiring additional data collection. For example, domain randomization [14, 15] typically alters environmental parameters like backgrounds in simulation or applies visual perturbations like color jitter. Generative models have also been used to augment object textures and add visual distractors [4, 16, 17]. However, explicit modeling and manipulation of lighting in real-world demonstrations remains underexplored, largely due to the difficulty of recovering geometric and material information required for simulating accurate light–material interactions.
@@ 1 Introduction
In this paper, we propose the first framework that applies physically-based inverse rendering for geometry- and material-aware lighting augmentation for imitation learning from real-world demonstrations in robotic manipulation. Our contributions are fourfold:
@@
  • Domain-Adapted Stable Video Diffusion. To ensure consistent augmentation across consecutive frames of each demonstration, we fine-tune Stable Video Diffusion (SVD) [19] for temporal lighting propagation. Fine-tuning data includes synthetic robot execution videos from Factor World [5] with varied lighting across tasks, and real-world execution videos from RoboNet [20] with visual degradation for radiometric accuracy.
@@
  • Real-world Evaluation. We validate our framework by assessing structural and temporal consistency in augmented demonstrations, and through real-world experiments on a 7-DoF robot with an embodiment unseen during SVD fine-tuning. Under six varied lighting conditions, our method reduces the behavior cloning generalization gap by 40.1% across 720 evaluations on two tasks, compared to models trained without lighting augmentation.
@@
  • Downstream Applications. We showcase augmentations on three additional environmental factors using geometric and material properties estimated by our framework.
2 Related Work
Data Augmentation for Robotic Manipulation.
@@ 2 Related Work
Domain randomization [21, 22, 23, 24] has been widely studied to enhance policy robustness when transferring from simulation to the real world. Building upon this, recent efforts [4, 17, 16, 25, 26] have focused on semantic augmentation of real-world images leveraging text-driven generative models [27, 28] to introduce texture variation, visual distractors, etc. For example, ROSIE [17] proposes changing object textures by first segmenting augmentation regions [29] and then performing text-guided image inpainting [30]. Rendering techniques [31, 32, 33] have also been applied for viewpoint augmentation [34, 35, 36, 37]. Another line of work breaks long-horizon tasks into object-centric subtasks or manipulation skills and replays transformed demonstrations in simulation to generate new data from a limited number of examples [38, 39, 40, 41, 42]. In this paper, we focus on the underexplored problem of lighting augmentation on real-world human demonstrations for robotic manipulation.
Inverse Rendering and Relighting.
@@
Relighting a scene typically requires identifying and altering its properties to produce the intended lighting effect [43]. Inverse rendering facilitates this by providing separate or joint estimations of geometry [44, 45, 46], material [47, 48, 49], and lighting [50, 51, 52] in a scene. Based on input requirements, inverse rendering can be categorized into single-view [53, 54, 55] and multi-view [56, 57, 58] methods. While methods such as DPI [59] and FIPT [60] produce high-fidelity relighting results, they rely on multi-view inputs for scene reconstruction and domain-specific datasets [61], making them incompatible with most existing robotic dataset camera setups. In this work, we present the first integration of single-view inverse rendering into robotic manipulation for explicit modeling of scene geometry and material properties.
3 Approach
3.1 Formulation and Overview
Given one episode of real-world human demonstration ($\mathbf{I}$, $\mathbf{P}$, $\mathbf{A}$) recorded from a fixed viewpoint, where $\mathbf{I}\in\mathbb{R}^{H\times W\times 3\times T}$ represents a sequence of $T$ RGB images, $\mathbf{P}\in\mathbb{R}^{p\times T}$ denotes proprioception states, and $\mathbf{A}\in\mathbb{R}^{a\times T}$ corresponds to actions, our goal is to perform physically-based lighting augmentation on $\mathbf{I}$ to produce a relit sequence $\mathbf{I}^{*}$ from the same viewpoint, while preserving structural and temporal consistency. Here, physically-based refers to physically-based rendering (PBR) for accurate light-material interactions.
Overview.
As shown in Fig. 2 the proposed framework consists of three components: (1) we begin by selecting the first frame $I_{0}$ from $\mathbf{I}$ and applying inverse rendering (Sec. 3.2) to estimate the geometric and material properties of the scene depicted in $I_{0}$; (2) the rendering module (Sec. 3.3) then uses the estimated properties and an environment map $E$ to relight $I_{0}$, producing $I_{0}^{*}$; and (3) the temporal propagation module (Sec. 3.4) propagates the lighting from $I_{0}^{*}$ across the full image sequence $\mathbf{I}$, resulting in the final augmented sequence $\mathbf{I}^{*}$. This forms an augmented episode ($\mathbf{I}^{*}$, $\mathbf{P}$, $\mathbf{A}$), which is then used to train the behavior cloning policy. The environment map $E$ has two use cases: (1) approximating the current lighting conditions to train a policy adapted to the environment (Sec. 4.2), and (2) introducing diverse lighting to contribute to a lighting-invariant policy (Sec. 5).
@@ 3.1 Formulation and Overview
Figure 2: Given the first frame $I_{0}$ in a real-world human demonstration, the inverse rendering module estimates material ($A^{*}$, $R^{*}$, $M^{*}$) and geometric ($N^{*}$, $D^{*}$) properties. The rendering module uses these estimates and an environment map $E$ providing new lighting to produce a relit frame $I_{0}^{*}$. The new lighting is then propagated across the entire sequence $\mathbf{I}$, producing $\mathbf{I}^{*}$ and forming an augmented episode ($\mathbf{I}^{*}$, $\mathbf{P}$, $\mathbf{A}$) for behavior cloning.
3.2 Single-Frame Inverse Rendering
@@
Following [18], we use a pre-trained network $\mathcal{P}$ to predict geometric and material properties of the scene depicted in $I_{0}$:
@@
where $I_{p}$ denotes the frame rendered using the predicted properties. $\mathcal{L}_{p}=\mathcal{L}_{\textit{re}}+\delta\mathcal{L}_{\textit{cons}}$, where $\mathcal{L}_{\textit{re}}$ denotes the reconstruction loss between the rendered frame $I_{p}$ and the original frame $I_{0}$, and $\mathcal{L}_{\textit{cons}}$ enforces consistency between the optimized and originally predicted properties, scaled by a factor $\delta$. Further details are provided in the supplementary material.
3.3 Single-Frame Relighting
@@
Light Sampling.
Given an environment map $E\in\mathbb{R}^{3}$, the incoming radiance $\lambda$ in direction $\omega_{i}$ is sampled as:
@@
where $E$ is obtained either by approximating [18] or measuring [62, 63] the current lighting condition using graphics techniques, or by selecting from open-source HDRI datasets [64, 65].
Scene Modeling.
The estimated properties from Sec. 3.2 are used to model the material appearance $\mathcal{M}$ along direction $\omega_{i}$ of the scene depicted in $I_{0}$, using the widely adopted Disney BRDF [66]:
@@
where $f_{\textit{diffuse}}$ and $f_{\textit{specular}}$ model the diffuse and specular components of reflection, respectively; further details are provided in the supplementary material.
Scene Relighting.
Using the simplified rendering equation [67], the augmented frame $I_{0}^{*}$ under new lighting from environment map $E$ is computed as:
@@
where $\Omega$ is the hemisphere centered at surface normal $N^{*}$, containing all incoming directions $\omega_{i}$. Note that once the property estimation for $I_{0}$ from Sec. 3.2 is completed, multiple augmentations can be applied using the same set of estimated properties.
3.4 Temporal Propagation
Domain-Adapted Stable Video Diffusion.
To meet the temporal consistency requirements of robotic manipulation, we adopt Stable Video Diffusion (SVD) [19] for its temporal modeling capabilities to propagate lighting from $I_{0}^{*}$ across the entire demonstration. We formulate this as a video-to-video translation task: lighting from the original image sequence $\mathbf{I}$ is transferred to generate the augmented sequence $\mathbf{I}^{*}$, guided by the reference frame $I_{0}^{*}$. This is defined as:
@@ 3.4 Temporal Propagation
Robotic Relighting Data Curation.
@@
Figure 3: (a) Examples of synthetic robot execution videos with lighting variations. (b) Examples of real-world robot execution videos with visual degradation.
@@
  • Synthetic videos with lighting variation ($\mathcal{D}_{1}$, Fig. 3(a)). We generate domain-specific relighting data for robotic manipulation using the Factor World [5] benchmark, introducing lighting variation under identical task execution trajectories. All 42 built-in scenes covering 19 tasks (e.g., door opening) are used, each captured with six camera views: one fixed top-down and five randomized within constrained azimuth, inclination, and radius in the range $[-\pi/2,\pi/2]$ to ensure robot visibility. Each scene includes 30 lighting conditions by sampling ambient and diffuse RGB values from [25,255][25,255][ 25 , 255 ]. Videos are rendered at $\times 512$ 512 resolution with 100 frames at 24 fps. For fine-tuning, we randomly select pairs of videos showing the same task trajectories under different lighting, using one as the input $\mathbf{I}$ and the other as the target sequence $\mathbf{I}^{*}$.
@@
  • Real-world videos with visual degradation ($\mathcal{D}_{2}$, Fig. 3(b)). Since existing real-world robotic datasets lack perfectly paired original and relit videos under identical task execution trajectories, we manually create video pairs by applying visual degradation on original videos from RoboNet [20] with 15 million video frames from 7 different robot platforms. We apply random visual transformations by sampling brightness, contrast, and saturation scaling factors from [0.2,1.9][0.2,1.9][ 0.2 , 1.9 ], and hue shifts from [0.5,0.5][-0.5,0.5][ - 0.5 , 0.5 ]. The degraded robot execution video serves as the input $\mathbf{I}$, while the original video acts as the output $\mathbf{I}^{*}$.
Fine-tuning.
@@
We fine-tune SVD in two stages. In Stage 1, we use synthetic videos from $\mathcal{D}_{1}$, which provide the main data for domain adaptation and reduce overfitting via controlled lighting variation. However, due to the non-physically-based lighting in Factor World, this stage lacks radiometric accuracy, motivating fine-tuning on real-world videos from $\mathcal{D}_{2}$ in Stage 2. Fine-tuning is performed on a pretrained 14-frame SVD model using a single H100 GPU for 8,000 steps for Stage 1 and 1,000 steps for Stage 2. A limited number of steps are applied to Stage 2 due to the absence of ground-truth relit pairs. Further details and ablation studies are provided in the supplementary material.
4 Experiments
4.1 Visual Lighting Quality Evaluation

### v2

# Physically-Based Lighting Generation for Robotic Manipulation

@@ Abstract
In this paper, we propose the first framework that leverages physically-based inverse rendering for novel lighting generation on existing real-world human demonstrations of robotic manipulation tasks. Specifically, inverse rendering decomposes the first frame in each demonstration into geometric (surface normal, depth) and material (albedo, roughness, metallic) properties, which are then used to render appearance changes under different lighting sources. To improve efficiency and maintain consistency across each generated sequence, we fine-tune Stable Video Diffusion on robot execution videos for temporal lighting propagation. We evaluate our framework by measuring the visual quality of the generated sequences, assessing its effectiveness in improving the imitation learning policy performance (38.75%) under six unseen real-world lighting conditions, and conduct ablation studies on individual modules of the proposed framework. We further showcase three downstream applications enabled by the proposed framework: background generation, object texture generation and distractor positioning. The code for the framework will be made publicly available.
I Introduction
Imitation learning from large-scale human demonstrations has proven to be an effective approach to deploying robotic manipulation tasks [1]. Yet collecting such data is costly, as it often needs to cover both diverse motor skills and varied visual appearances [2, 3]. In a fixed object-environment setup, achieving reliable policy performance on a single skill typically requires a proficient operator to repeat the task about 200 times using specialized teleoperation devices [4]. To ensure broader robustness, dataset construction further incorporates visual variations such as object texture, background, and distractors [5, 6]. Capturing these variations requires repeating the data collection process for each factor, making it especially costly to achieve sufficient visual coverage for every skill in real-world settings.
In response, substantial effort has been devoted to synthesizing object texture and background variations [2, 7, 8, 9, 10]. Inpainting techniques based on generative modeling [8] and physical tools such as greenscreens [9] are employed, yielding promising results while reducing the need for additional data collection. By comparison, lighting, another pervasive and highly dynamic factor in real-world settings, has received little attention. Even in the relatively controlled indoor environments, lighting can vary significantly with artificial sources such as lamps or daylight through windows. In outdoor environments, the variability is far greater, as field robots may be exposed to natural lighting changes such as those caused by the position of the sun or changes in cloud cover. Such variability points to the need for training data under different lighting. However, given the existing cost of real-world data collection, expanding datasets to cover each skill under lighting variations only compounds the challenge.
Meanwhile, lighting poses a unique modeling challenge. As light travels from its source, it scatters and reflects off all scene components before ultimately being captured by a camera. Consequently, variations in lighting alter the appearance of the entire scene and influence the training process that depends on camera observations. This recursive property of lighting makes it extremely difficult to synthesize in robotic scenes. One example is shown in Fig. 1(a): when a lamp is turned off, all scene components appear darker (global impact), while nearby components alter one another’s appearance by casting shadows (local impact). Such behaviors may also help explain why many studies report that trained policies are highly vulnerable to lighting variations [3, 6, 11]. Taken together, these bring us to the question:
Can we generate lighting that approximates real-world variations to reduce costly data collection for robotic manipulation?
Real
Generated
Figure 1: (a) Example of recursive effect of lighting: when a lamp is turned off, all scene components appear darker (global impact), while nearby components alter one another’s appearance by casting shadows (local impact). (b) Left: existing real-world human demonstration. Right: demonstration relit by our method. Our framework generates novel lighting for real-world demonstrations to approximate scenes under unseen lighting conditions.
@@ I Introduction
To tackle this, we propose RoLight, the first framework that leverages physically-based inverse rendering for robotic scene lighting generation. Inverse rendering is introduced for explicit modeling of geometric and material information in existing real-world human demonstrations for simulating accurate light-material interactions. Our contributions are fourfold:
@@
  • Domain-Adapted Stable Video Diffusion. To improve efficiency and maintain consistent generation across consecutive frames of each demonstration, we fine-tune Stable Video Diffusion (SVD) [13] for temporal lighting propagation. Fine-tuning data includes synthetic robot execution videos from Factor World [3] with varied lighting across tasks, and real-world execution videos from RoboNet [14] with visual degradation for radiometric accuracy.
@@
  • Real-world Evaluation. We validate our framework by assessing structural and temporal consistency in generated demonstrations, and through real-world experiments on a 7-DoF robot with an embodiment unseen during SVD fine-tuning. Under six varied lighting conditions, our method improves the imitation learning policy performance by 38.75% across 1,000 evaluations on two tasks, compared to models trained without lighting generation. Ablation studies are further conducted on individual modules of the proposed framework.
@@
  • Downstream Applications. We showcase generations on three additional environmental factors using geometric and material properties estimated by our framework.
II Related Work
II-A Data Generation for Robotic Manipulation
@@
To address the robotic data bottleneck, recent efforts [2, 7, 8, 15] have focused on semantic augmentation of real-world images leveraging text-driven generative models [16] to introduce texture variation, visual distractors, etc. For example, ROSIE [2] proposes changing object textures by first segmenting generation regions [17] and then performing text-guided image inpainting [18]. Rendering techniques [19] have also been applied for viewpoint generation [20, 21, 22]. Another line of work breaks long-horizon tasks into object-centric subtasks or manipulation skills and replays transformed demonstrations in simulation to generate new data from a limited number of examples [4, 23, 24]. In this paper, we focus on the underexplored problem of novel lighting generation on real-world human demonstrations for robotic manipulation.
II-B Inverse Rendering and Relighting
@@
Relighting a scene typically requires identifying and altering its properties to produce the intended lighting effect [25]. Inverse rendering facilitates this by providing separate or joint estimations of geometry [26], material [27], and lighting [28] in a scene. Based on input requirements, inverse rendering can be categorized into single-view [29] and multi-view [30] methods. While methods such as DPI [31] and FIPT [32] produce high-fidelity relighting results, they rely on multi-view inputs for scene reconstruction and domain-specific datasets [33], making them incompatible with most existing robotic dataset camera setups. In this work, we present the first integration of single-view inverse rendering into robotic manipulation for explicit modeling of scene geometry and material properties.
II-C Latent Video Diffusion Models
Latent diffusion models [16] generate images by iteratively denoising Gaussian noise to approximate the target distribution. Latent video diffusion models [13, 34] extend this framework to the video domain typically by introducing temporal mixing layers on top of pre-trained image generation architectures. For example, Stable Video Diffusion [13] extends Stable Diffusion [16] by inserting temporal convolution and attention layers after each spatial layer and finetuning on curated video data. The extended temporal coherence combined with flexible conditioning (e.g., text, reference frames) enables applications such as camera-controlled video generation [35], multi-view synthesis [36], and video prediction [37]. In this work, we adopt this structure for video-to-video translation, transferring lighting in real-world human demonstrations to new lighting conditions using physically-based relit frames as reference.
@@ II-C Latent Video Diffusion Models
Figure 2: Given the first frame $I_{0}$ in a real-world human demonstration, the inverse rendering module estimates material ($A^{*}$, $R^{*}$, $M^{*}$) and geometric ($N^{*}$, $D^{*}$) properties. The rendering module uses these estimates and an environment map $E$ providing new lighting to produce a relit frame $I_{0}^{*}$. The new lighting is then propagated across the entire sequence $\mathbf{I}$, producing $\mathbf{I}^{*}$ and forming an generated episode ($\mathbf{I}^{*}$, $\mathbf{P}$, $\mathbf{A}$) for behavior cloning.
III Methodology
III-A Formulation and Overview
Given one episode of real-world human demonstration ($\mathbf{I}$, $\mathbf{P}$, $\mathbf{A}$) recorded from a fixed viewpoint, where $\mathbf{I}\in\mathbb{R}^{H\times W\times 3\times T}$ represents a sequence of $T$ RGB images, $\mathbf{P}\in\mathbb{R}^{p\times T}$ denotes proprioception states, and $\mathbf{A}\in\mathbb{R}^{a\times T}$ corresponds to actions. Our goal is to transform $\mathbf{I}$ into $\mathbf{I}^{*}$ under new lighting, using physically based rendering (PBR) to ensure accurate light-material interactions.
Overview. As shown in Fig. 2 the proposed framework consists of three components: (1) we begin by selecting the first frame $I_{0}$ from $\mathbf{I}$ and applying inverse rendering (Sec. III-B) to estimate the geometric and material properties of the scene depicted in $I_{0}$; (2) the rendering module (Sec. III-C) then uses the estimated properties and an environment map $E$ to relight $I_{0}$, producing $I_{0}^{*}$; and (3) the temporal propagation module (Sec. III-D) propagates the lighting from $I_{0}^{*}$ across the full image sequence $\mathbf{I}$, resulting in the final generated sequence $\mathbf{I}^{*}$. This forms an generated episode ($\mathbf{I}^{*}$, $\mathbf{P}$, $\mathbf{A}$), which is then used to train the imitation learning policy. The environment map $E$ has two use cases: (1) approximating the current lighting conditions to train a policy adapted to the current environment (Sec. IV-B), and (2) introducing diverse lighting to contribute to a lighting-invariant policy (Sec. V).
III-B Single-Frame Inverse Rendering
@@
Following [12], we use the pre-trained network $\mathcal{P}$ to predict geometric and material properties of the scene depicted in $I_{0}$:
@@
where $I_{p}$ denotes the frame rendered using the predicted properties. $\mathcal{L}_{p}=\mathcal{L}_{\textit{re}}+\delta\mathcal{L}_{\textit{cons}}$, where $\mathcal{L}_{\textit{re}}$ denotes the reconstruction loss between the rendered frame $I_{p}$ and the original frame $I_{0}$, and $\mathcal{L}_{\textit{cons}}$ denotes the $L_{1}$ consistency loss the optimized and originally predicted properties, scaled by a factor $\delta$ set to 0.005 in most experiments.
III-C Single-Frame Relighting
@@
Target Light Sampling
Given an environment map $E\in\mathbb{R}^{3}$ containing target lighting condition, the incoming radiance $\lambda$ in direction $\omega_{i}$ is sampled as:
@@
where $E$ is obtained either from open-source HDRI liraries [38] (Fig. 3(a)), by optimization during inverse rendering [12] (Fig. 3(b)) or by measuring the current lighting condition using graphics techniques [39] (Fig. 3(c)).
Figure 3: (a) Environment maps from open-source HDRI libraries retrieved using keywords such as “blue studio” and “sunrise in the field”. (b) Environment maps derived via optimization with inverse rendering, where resolution depends on the specific method. (c) Environment maps obtained by photographing a chrome (mirror) ball under multiple exposures and unwrapping the result.
Scene Modeling
The estimated properties from Sec. III-B are used to model the material appearance $\mathcal{M}$ along direction $\omega_{i}$ of the scene depicted in $I_{0}$, using the widely adopted Disney BRDF [40]:
@@
where $f_{\textit{diffuse}}$ and $f_{\textit{specular}}$ represent the diffuse and specular reflection components of the Disney BRDF, respectively.
Scene Relighting
Using the simplified rendering equation [41], the generated frame $I_{0}^{*}$ under new lighting from environment map $E$ is computed as:
@@
where $\Omega$ is the hemisphere centered at surface normal $N^{*}$, containing all incoming directions $\omega_{i}$. Note that once the property estimation for $I_{0}$ from Sec. III-B is completed, multiple generations can be applied using the same set of estimated properties.
III-D Temporal Propagation
Domain-Adapted Stable Video Diffusion
To improve efficiency and maintain consistency in the generated sequences, we adopt Stable Video Diffusion (SVD) [13] for its temporal modeling capabilities to propagate lighting from $I_{0}^{*}$ across the entire demonstration. We formulate this as a video-to-video translation task: lighting from the original image sequence $\mathbf{I}$ is transferred to generate the new sequence $\mathbf{I}^{*}$, guided by the reference frame $I_{0}^{*}$. This is defined as:
@@ III-D Temporal Propagation
Robotic Relighting Data Curation
@@
  • Synthetic videos with lighting variation ($\mathcal{D}_{1}$, Fig. 4(a)). We generate domain-specific relighting data for robotic manipulation using the Factor World [3] benchmark, introducing lighting variation under identical task execution trajectories. All 42 built-in scenes covering 19 tasks (e.g., door opening) are used, each captured with six camera views: one fixed top-down and five randomized within constrained azimuth, inclination, and radius in the range $\pi/2,\pi/2]$ to ensure robot visibility. Each scene includes 30 lighting conditions by sampling ambient and diffuse RGB values from $[25,255]$. Videos are rendered at $512\times 512$ resolution with 100 frames at 24 fps. For fine-tuning, we randomly select pairs of videos showing the same task trajectories under different lighting, using one as the input $\mathbf{I}$ and the other as the target sequence $\mathbf{I}^{*}$.
@@
  • Real-world videos with visual degradation ($\mathcal{D}_{2}$, Fig. 4(b)). Since existing real-world robotic datasets lack perfectly paired original and relit videos under identical task execution trajectories, we manually create video pairs by applying visual degradation on original videos from RoboNet [14] with 15 million video frames from 7 different robot platforms. We apply random visual transformations by sampling brightness, contrast, and saturation scaling factors from $[0.2,1.9]$, and hue shifts from [0.5,0.5][-0.5,0.5]. The degraded robot execution video serves as the input $\mathbf{I}$, while the original video acts as the output $\mathbf{I}^{*}$.
Figure 4: (a) Examples of synthetic robot execution videos with lighting variations ($\mathcal{D}_{1}$). (b) Examples of real-world robot execution videos with visual degradation ($\mathcal{D}_{2}$).
Fine-tuning
@@
We fine-tune SVD in two stages. In Stage 1, we use synthetic videos from $\mathcal{D}_{1}$, which provide the main data for domain adaptation and reduce overfitting via controlled lighting variation. However, due to the non-physically-based lighting in Factor World, this stage lacks radiometric accuracy, motivating fine-tuning on real-world videos from $\mathcal{D}_{2}$ in Stage 2. Fine-tuning is performed on a pretrained 14-frame SVD model using a single H100 GPU for 8,000 steps for Stage 1 and 1,000 steps for Stage 2. A limited number of steps are applied to Stage 2 due to the absence of ground-truth relit pairs.
IV Experiments
IV-A Visual Lighting Quality Evaluation